• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Armstrong

    Log entry time 23:52:55 on July29,2003

    Entry number 108067

    keyword=Target Shift Log


    Target Log   Swing shift   July 29 2003



    D.S. Armstrong




    Shift Start:

    holding field: 0°
    target HWP: out
    lasers: 4,6,7 on
    Target: Pol 3He
    Cell: Priapus
    Beam: waiting for beam





       At start of shift, waiting for beam   (plus ça change...)

      We are at start of kinematics 5.17.

    16:30   MCC sends tune beam into the hall; looks good, ask
    for 1 µA so that we can run spot++ and check raster.

    16:36    Spot looks good, ask for 4 µA.

    17:23    MCC is working on the Fast Feedback system;
    move to empty target position for caution's sake.

    17:52    Fast Feedback work is completed, and they
    believe that it will now be working. Move back to
    3He cell.

    18:00    Sirish suggests that we check the intrinsic
    spot size. We move to empty target position in order to
    do harp scans, OTR...

    18:45    Harp scans and raster adjustment done; back to
    3He cell.

    19:15    Turn off lasers 4,6,7: start rotation from
    0° to 270°.

    19:20    Turn on Laser 1
    19:21    Turn on Laser 2
    19:22    Turn on Laser 3

    19:48 Take away beam to go to new kinematics. Turn off
    lasers 1,2,3 and start rotation to 0°.

    19:55 Rotated. Turn on Laser 4
    19:56          Turn on Laser 6
    19:57         Turn on Laser 7

          waiting for temperature to stabilize...

    20:07    NMR measurement. Result = 41.63%
    20:13    EPR measurement (by Vince). Result = 40.6%

      We are now at start of kinematics
    5.18
    .

    20:15    Move to empty reference cell. 10 minute run.
    20:38    Fill Reference cell with 140 psig N2. 10 minute run.

    20:56    Move to 3He cell.
          While it is moving, vent and evacuate Reference cell.

    21:03    Cell now in position. Ask for
    3 µA beam.

    22:52    ask for beam off. Turn off Lasers 4,6,7,
    and start rotation of holding field to 270°.

    22:55    Field Rotated. Turn on Laser 1
    22:56    Turn on Laser 2
    22:57    Turn on Laser 3; ask for beam back.

    23:30    Take beam away; Turn off transverse lasers
    and move to reference cell position. Rotate holding field back to


    23:40    Turn on Laser 4
    23:41    Turn on Laser 6
    23:34    Turn on Laser 7




    Shift End:

    holding field: 0 °
    target HWP: out
    lasers: longitudinal: 4,6,7 ON
    Target: empty Reference cell
    Cell: Priapus
    Beam: 10.0 µA



    hand off to owl shift, go home to sleep...